OpenStudy (anonymous):

What is the quotient: (56x3 + 16x2 – 32) ÷ 8 ? 7x3 + 2x2 – 4x 7x3 – 2x2 – 4x 7x3 + 2x2 – 4 7x3 – 2x2 – 4

OpenStudy (anonymous):

divide each term by 8 for example 56(x^(3))/8 = 7 x^3

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

you've got \[( 56x ^{3} + 16x ^{2} - 32) \div8\] so divide each term by 8: \[\frac{56x ^{3} }{ 8 }\]= \[7x^{3}\], and do the others (16x^(2)/8 = 2x^2); (32/8 = 4) and you will get the final equation like 7x^3 + 2x^2 - 4

OpenStudy (anonymous):

thanks!!